How would I do this problem? Express 5 over 3 - radical 2 with a rational denominator, in simplest radical form.
Answer by edjones(8007)

5/(3-sqrt(2))
=5/(3-sqrt(2)) * (3+sqrt(2))/(3+sqrt(2))
=5(3+sqrt(2))/(9-2)
= 5(3+sqrt(2))/7
.
Ed